A valószínűségszámítás és matematika statisztika gyakorlaton való részvétel

dátum:  IX. 12. IX. 19. IX. 26. X. 3. X. 10. X. 17. X. 24.

diákok:

Araczki György  +  +  +  +
Balmaz Tamás  +  +  +  +  +
Bardóczi Ákos  +  +  +
Benke Zsolt  +  −  −  +  +
Bónizs Attila  +  +  −  +  +
Csernai Csaba  +  +  −  +
Dubravszky Csaba  +  +  +  +
Fábián Máté  +  +  +  +
Fülöp Gábor  +  +  +  +
Gali Gyula  +  +  +  −
Gilányi Ádám  +  −  −  +  +
Göncz Gábor  +  −  +  +  −
Hegedűs Péter  +  −  −  −
Helmeczi József  +  +  −  +
Hudák Levente  +  −
Karakas Péter  +  +  +  +  +
Kasza Imre Ádám  +  +  −  +  +
Keszeg Alpár  +  +  +  +  +
Kovács László  +  −  −  −  −  −
Kovács Lilla  +  +  −  +  +
Kovács Zoltán  +  +  +  +  +
Kozma Zoltán  +  +  −  +  +
Kulcsár István  +  −  −  −  −
Mező Péter  +  +  −
MD Monirul Babar Duke  +  +  +  +  +
Miskolczi Zsolt  +  −  +  +  +
Nagy István  +  −  −  −
Nyilas Attila  +  +  +  +  +
Nyúzó Tamás  +  +  −  −  +
Sikolya Tibor  +
Szabó Ernő  +  +  +  +  +  +
Terdik Sándor  +  +  +  +  +
Varga Tamás  +  −
Veres Gergő Márton  +  −  +
Vernika Tamás  +  +  +  +  +  +

dátum:  X. 31. XI. 7. XI. 14. XI. 21. XI. 28. XII. 5. XII. 12.

diákok:

Araczki György  −  −  +  −  −  −
Balmaz Tamás  +  +  +  +  +  +  +
Bardóczi Ákos  −  +  +  −  −  +  +
Benke Zsolt  +  −  −  −  −  −
Bónizs Attila  +  +  +  +  +  +  +
Csernai Csaba  +  −  +  +  −  −
Dubravszky Csaba  −  +  +  +  +  +  +
Fábián Máté  +  −  −  +  +  +  +
Fülöp Gábor  −  −  −  +  −  −
Gilányi Ádám  +  −  −  −  −  −
Göncz Gábor  +  +  −  +  −  −  +
Karakas Péter  +  +  +  −  +  +  +
Kasza Imre Ádám  +  −  +  −  +  +  +
Keszeg Alpár  +  +  −  +  −  −
Kovács Lilla  +  +  +  +  −  +
Kovács Zoltán  −  −  +  +  +  −  +
Kozma Zoltán  +  +  +  +  +  −  +
MD Monirul Babar Duke  +  +  +  +  +  −
Miskolczi Zsolt  +  +  +  +  +  −  +
Nyilas Attila  +  +  +  +  −  +  +
Nyúzó Tamás  +  +  −  −  −  −
Szabó Ernő  −  +  +  +  +  +  +
Terdik Sándor  +  −  +  +  +  +
Veres Gergő Márton  +  +  −  +  −  −
Vernika Tamás  +  +  +  +  +  −  +